Create a sample of ten numbers that have a mean of 3000 psi (compressive strength of concrete). How did you coming up with these ten numbers demonstrate the concept of N-1 degrees of freedom that is used when calculating the standard deviation of a sample?
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 3000

sample standard deviation = s

sample size = n = 10

Degrees of freedom = df = n - 1 = 10 - 1 = 9

=

WE need to find sample standard deviation using this formula,

s2 =  (x - )2 / n - 1